OpenStudy (anonymous):

Forgot What to do here, HELP. Find the radius of a circle with the given circumference. 6√2 in.

OpenStudy (anonymous):

circumference = 2 * pi * radius. circumference = \[6\sqrt{2}\] to get radius, divide 6(2)^1/2 by 2 pi and get the answer.